# The Seesaw Evaded Modular Dirac Framework

### v2

@@ Introduction.
In such scenarios, the smallness of neutrino masses can be achieved via the Dirac seesaw mechanism [5], analogous to the Majorana case. While well motivated, these constructions often require additional symmetries to forbid tree-level Yukawa couplings and remain structurally similar to the conventional Majorana seesaw.
Motivated by these challenges, we argue that small neutrino masses can also be generated in a Dirac framework without invoking any seesaw mechanism. However, in such non seesaw scenarios, the burden of producing sub eV neutrino masses falls entirely on the couplings, which then must be unnaturally small.
@@
Interestingly, in one of our earlier works [6], we constructed a Dirac model without seesaw in which tiny neutrino masses emerge naturally from the interplay of flavon vevs and the cutoff scale, thereby avoiding both the seesaw mechanism and unnaturally suppressed couplings. While successful, that setup required an extended scalar sector, reducing its minimality.
@@
A natural way to overcome this limitation is to leverage the power of modular symmetry [7, 8, 9, 10], where both matter fields and Yukawa couplings transform under a finite modular subgroup $\Gamma_{N}$, with Yukawa couplings represented by modular forms. This framework substantially reduces the need for flavon fields and provides a minimal and elegant approach to understanding lepton masses and mixing. Although most studies implement modular symmetry in a supersymmetric (SUSY) context, low energy SUSY remains experimentally elusive, motivating the exploration of non SUSY modular frameworks. Recently ‘Qu’ and ‘Ding’ showed that modular symmetries can be consistently formulated in a non holomorphic framework [11], enabling their application in non SUSY models [12, 13, 14, 15, 16, 17, 18, 19, 20]. This non holomorphic modular approach therefore offers a minimal and predictive framework for flavour model building, while simultaneously reducing the reliance on an extended scalar sector.
@@
In this work, we adopt a novel approach by presenting a non SUSY modular framework, based on the finite modular subgroup $\Gamma_{3}\simeq A_{4}$, in which neutrinos are purely Dirac particles, independent of the seesaw mechanism, so that neutrino masses arise naturally without requiring unnaturally suppressed couplings.
Model setup.
@@
We extend the SM by incorporating the modular $A_{4}$ flavour symmetry [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The field content relevant for the lepton sector is as follows: the lepton doublets $\overline{L}_{l}$ transform as a triplet under $A_{4}$ with weight $k=1$, the right handed neutrinos $\nu_{R}$ transform as a triplet under $A_{4}$ with weight 1-1, and the right handed charged leptons $(e_{R},\mu_{R},\tau_{R})$ transform as $1,1^{\prime\prime},1^{\prime}$ under $A_{4}$ with weights $5$, 33, and $2$, respectively. The Higgs doublet $H$ is an $A_{4}$ singlet with weight 0, and the singlet scalars $\rho,\sigma,\eta$ are all $A_{4}$ singlets with weights +2,1,+2,-1, and +1+1, respectively. In our construction, only modular forms of weights +2+2 and +4+4 are used, ensuring that Majorana mass terms for $\nu_{R}$ are absent and neutrinos remain Dirac. However, a modular form of weight 2-2 could, in principle, reintroduce Majorana operators. While it is common in modular models to assume that certain modular weights do not appear in a theory. In our model, this would correspond to weight 2-2. However, such an assumption is not fully justified. To ensure all Majorana terms are forbidden, we instead impose an additional $Z_{10}$ symmetry [33, 6, 34, 35]. Under this symmetry, only $\overline{L}_{l}$, the right handed charged leptons, and $\nu_{R}$ transform non trivially, with charges $6$, 44, and $4$, respectively. These assignments consistently forbid all Majorana operators, maintaining a purely Dirac neutrino framework.
@@
$\sin^{2}\theta_{23}$ $=\frac{|U_{23}|^{2}}{1-|U_{13}|^{2}},\qquad\delta=-\text{Arg}[U_{13}].$ (5)
We fix $v_{h}=246~\text{GeV}$, $\Lambda=10^{14}~\text{GeV}$, $v_{\sigma}=10^{12}~\text{GeV}$, $\beta_{\mu}/\alpha_{e}=0.0203$, and $\gamma_{\tau}/\alpha_{e}=0.003$, and perform a numerical scan over the remaining parameters, including the modulus $\tau$ restricted to its fundamental domain [8]. The objective is to identify regions of parameter space that yield a hierarchical charged lepton spectrum while simultaneously reproducing the mixing angles and the Dirac CP violating phase, and the mass squared differences $(\Delta m_{21}^{2},\Delta m_{31}^{2})$ at their global (NuFIT 6.0) best fit values [36]. In addition, the sum of neutrino masses $\sum m_{i}$ is required to remain consistent with cosmological bound [37]. Although other regions of the parameter space can reproduce the observables within their $3\sigma$ ranges, our focus here is on identifying the regions of parameter space that are in excellent agreement with the global best fit values, as these correspond to the most probable configuration. The scan is carried out for both normal hierarchy (NH) and inverted hierarchy (IH).
@@
$v_{\eta}/10^{12}$ $8.752~\text{GeV}$ $9.694~\text{GeV}$ $16~\text{GeV}$ $8.147~\text{GeV}$
@@
From Table 1, it is evident that, for both NH and IH, the allowed parameter space is tightly constrained by the best fit requirement. The modulus $\tau$ lies within a narrow band of the fundamental domain, with $Im[\tau]$ clustering around 1.01.51.0-1.5; for these values of $\tau$, the neutrino mixing angles and mass squared differences are reproduced in excellent agreement with their global best fit values. In the NH case, while the CP phase $\delta$ does not exactly coincide with the best fit point, the prediction $\delta\approx 270^{\circ}$ remains consistent with the NuFIT 6.0 global analysis, favouring maximal CP violation. In the charged lepton sector, the observed mass hierarchy is realized, with $m_{e}\simeq 0.5~\text{MeV}$, $m_{\mu}\simeq 105~\text{MeV}$, and $m_{\tau}\simeq 1777~\text{MeV}$. The framework also predicts a sharply determined sum of neutrino masses, $\Sigma m_{i}\simeq 0.10$ eV (NH) and $\Sigma m_{i}\simeq 0.115$ eV (IH), which is consistent with current cosmological bound. As noted previously, owing to the rich structure of the model, a broader scan of the parameter space can reproduce other regions of the $3\sigma$ ranges of the observables. Consequently, even if future global fits deviate from the current best fit values, the model can be tuned within this extended parameter space to maintain consistency, thereby preserving its predictive power and phenomenological viability. In addition, the scalar vevs are driven into distinct ranges: $v_{\eta}\sim\mathcal{O}(10^{12})~\text{GeV}$ and $v_{\rho}$ in the electroweak – TeV window (30030010001000 GeV). The parameters $y_{i}$ are found to take reasonable values compatible with the charged lepton mass hierarchy and the neutrino sector observables.
@@
Summary.
We present a novel, seesaw evaded Dirac neutrino framework based on modular $A_{4}$ symmetry. In this framework, sub eV neutrino masses arise naturally from the interplay of modular forms, three singlet scalar vevs, and the cutoff scale, without requiring unnaturally small couplings. The model uniquely predicts all neutrino mixing angles and mass squared differences consistent with the global best fit values. In addition, it yields a sum of neutrino masses compatible with cosmological bounds and predicts maximal Dirac CP violation. Owing to its economy of fields and predictive character, this framework offers a fresh perspective on the origin of neutrino masses and mixing beyond traditional discrete flavour symmetry and seesaw based approaches.
@@ Acknowledgements.
MD is thankful to Subhankar Roy, Department of Physics, Gauhati University, for his kind support and encouragement. MD also acknowledges the Council of Scientific and Industrial Research (CSIR), Government of India, New Delhi, for financial support through a NET Senior Research Fellowship, vide grant No. 09/0059(15346)/2022-EMR-I.

### v3

@@ Abstract
In such scenarios, the smallness of neutrino masses can be achieved via the Dirac seesaw mechanism [5], analogous to the Majorana case. While well motivated, we argue that small neutrino masses can also be generated in a Dirac framework without invoking any seesaw mechanism. However, in such non seesaw scenarios, the renormalisable tree level term is generally allowed, and reproducing sub eV neutrino masses then requires the tree level Yukawa couplings to be unnaturally small. Therefore, any realistic non seesaw Dirac model must first forbid the tree level Dirac Yukawa and then generate neutrino masses through a suitably suppressed higher dimensional operator.
@@
Interestingly, in one of our earlier works [6], we constructed a Dirac model without a seesaw mechanism, in which the tree level term was forbidden by discrete symmetries. In that setup, tiny neutrino masses emerged naturally from the interplay of flavon vevs and the cutoff scale, thereby avoiding both the seesaw mechanism and unnaturally small Yukawa couplings. While successful, that construction required an extended scalar sector, reducing its minimality.
@@
A natural way to overcome this limitation is to leverage the power of modular symmetry [7, 8, 9, 10], where both matter fields and Yukawa couplings transform under a finite modular subgroup $\Gamma_{N}$, with Yukawa couplings represented by modular forms. This framework substantially reduces the need for flavon fields and provides a minimal and elegant approach to understanding lepton masses and mixing. Although most studies implement modular symmetry in a supersymmetric (SUSY) context, low energy SUSY remains experimentally elusive, motivating the exploration of non SUSY modular frameworks. Recently ‘Qu’ and ‘Ding’ showed that modular symmetries can be consistently formulated in a non holomorphic framework [11], enabling their application in non SUSY models [12, 13, 14, 15, 16, 17, 18, 19, 20]. This non holomorphic modular approach therefore offers a predictive framework for flavour model building, while simultaneously reducing the reliance on an extended scalar sector.
@@
In this work, we adopt a novel approach by presenting a non SUSY modular framework, based on the finite modular subgroup $\Gamma_{3}\simeq A_{4}$, in which neutrinos are purely Dirac particles, independent of the seesaw mechanism. Although seesaw realizations have been the dominant route in modular symmetry models, our framework demonstrates that neutrino masses can arise naturally without invoking a seesaw and without requiring unnaturally suppressed couplings.
@@
We extend the SM by incorporating the modular $A_{4}$ flavour symmetry [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The field content relevant for the lepton sector is as follows: the lepton doublets $\overline{L}_{l}$ transform as a triplet under $A_{4}$ with weight $k=1$, the right handed neutrinos $\nu_{R}$ transform as a triplet under $A_{4}$ with weight 1-1, and the right handed charged leptons $(e_{R},\mu_{R},\tau_{R})$ transform as $1,1^{\prime\prime},1^{\prime}$ under $A_{4}$ with weights $5$, 33, and $2$, respectively. The Higgs doublet $H$ is an $A_{4}$ singlet with weight 0, and the singlet scalars $\rho,\sigma,\eta$ are all $A_{4}$ singlets with weights +2,1,+2,-1, and +1+1, respectively. In our construction, only modular forms of weights +2+2 and +4+4 are used, ensuring that Majorana mass terms for $\nu_{R}$ are absent and neutrinos remain Dirac. However, a modular form of weight 2-2 could, in principle, reintroduce Majorana operators. While it is common in modular models to assume that certain modular weights do not appear in a theory. In our model, this would correspond to weight 2-2. However, such an assumption is not fully justified. To ensure all Majorana terms are forbidden, we instead impose an additional $Z_{10}$ symmetry [32, 6, 33, 34]. Under this symmetry, only $\overline{L}_{l}$, the right handed charged leptons, and $\nu_{R}$ transform non trivially, with charges $6$, 44, and $4$, respectively. These assignments consistently forbid all Majorana operators, maintaining a purely Dirac neutrino framework.
@@
$\sin^{2}\theta_{23}=\frac{|U_{23}|^{2}}{1-|U_{13}|^{2}},\qquad\sin\delta=\frac{J_{\text{CP}}}{c_{12}s_{12}c_{23}s_{23}c_{13}^{2}s_{13}},$ (5)
where, $J_{\text{CP}}=\operatorname{Im}\left[U_{11}U_{22}U_{12}^{*}U_{21}^{*}\right]$ is the Jarlskog invariant, and $c_{ij}=\cos\theta_{ij}$, $s_{ij}=\sin\theta_{ij}$. For the numerical analysis, we fix $v_{h}=246~\text{GeV}$, $\Lambda=10^{14}~\text{GeV}$, $v_{\sigma}=10^{12}~\text{GeV}$, $\beta_{\mu}/\alpha_{e}=0.0203$, and $\gamma_{\tau}/\alpha_{e}=0.003$, and perform a numerical scan over the remaining parameters, including the modulus $\tau$ restricted to its fundamental domain [8]. The objective is to identify regions of parameter space that yield a hierarchical charged lepton spectrum while simultaneously reproducing the mixing angles and the Dirac CP violating phase, and the mass squared differences $(\Delta m_{21}^{2},\Delta m_{31}^{2})$ at their global (NuFIT 6.0) best fit values [35]. In addition, the sum of neutrino masses $\sum m_{i}$ is required to remain consistent with cosmological bound [36]. Although other regions of the parameter space can reproduce the observables within their $3\sigma$ ranges, our focus here is on identifying the regions of parameter space that are in excellent agreement with the global best fit values, as these correspond to the most probable configuration. The scan is carried out for both normal hierarchy (NH) and inverted hierarchy (IH).
@@
$|v_{\eta}|/10^{12}$ $8.752~\text{GeV}$ $9.350~\text{GeV}$ $8.147~\text{GeV}$ $16~\text{GeV}$
@@
From Table 1, it is evident that, for both NH and IH, the allowed parameter space is tightly constrained by the best fit requirement. The modulus $\tau$ lies within a narrow band of the fundamental domain, with $Im[\tau]$ clustering around 1.01.51.0-1.5; for these values of $\tau$, the neutrino mixing angles and mass squared differences are reproduced in excellent agreement with their global best fit values. In the NH case, while the CP phase $\delta$ does not exactly coincide with the best fit point, the prediction $\delta\approx 270^{\circ}$ remains consistent with the NuFIT 6.0 global analysis [35], favouring maximal CP violation. In the charged lepton sector, the observed mass hierarchy is realized, with $m_{e}\simeq 0.5~\text{MeV}$, $m_{\mu}\simeq 105~\text{MeV}$, and $m_{\tau}\simeq 1777~\text{MeV}$. The framework also predicts a sharply determined sum of neutrino masses, $\Sigma m_{i}\simeq 0.10$ eV (NH) and $\Sigma m_{i}\simeq 0.115$ eV (IH), which is consistent with current cosmological upper bound [36]. As noted previously, owing to the rich structure of the model, a broader scan of the parameter space can reproduce other regions of the $3\sigma$ ranges of the observables. Consequently, even if future global fits deviate from the current best fit values, the model can be tuned within this extended parameter space to maintain consistency, thereby preserving its predictive power and phenomenological viability. In addition, the scalar vevs are driven into distinct ranges: $v_{\eta}\sim\mathcal{O}(10^{12})~\text{GeV}$ and $v_{\rho}$ in the electroweak – TeV window (30030010001000 GeV). The parameters $y_{i}$ are found to take reasonable values compatible with the charged lepton mass hierarchy and the neutrino sector observables.
@@
In sum, we present a novel, seesaw evaded Dirac neutrino framework based on modular $A_{4}$ symmetry. In this framework, sub eV neutrino masses arise naturally from the interplay of modular forms, three singlet scalar vevs, and the cutoff scale, without requiring unnaturally small couplings. The model uniquely predicts all neutrino mixing angles and mass squared differences consistent with the global best fit values. In addition, it yields a sum of neutrino masses compatible with cosmological bounds and predicts maximal Dirac CP violation. Owing to its economy of fields and predictive character, this framework offers a fresh perspective on the origin of neutrino masses and mixing beyond traditional discrete flavour symmetry and seesaw based approaches.
@@ Acknowledgements.
MD is thankful to Subhankar Roy, Department of Physics, Gauhati University, for his kind support and encouragement. MD would also like to thank Amol Dighe (TIFR, Mumbai) and S. Uma Sankar (IIT Bombay) for valuable discussions and comments on this work during the IIFC-vP Neutrino School at NISER, Bhubaneswar, 2025. MD also acknowledges the Council of Scientific and Industrial Research (CSIR), Government of India, New Delhi, for financial support through a NET Senior Research Fellowship, vide grant No. 09/0059(15346)/2022-EMR-I, under which this work was carried out.